In a population of gorillas scientists sample the tyrosinase gene (tyr).  The tyr+ allele results in normal pigmentation in both heterozygotes and homozygotes.  The tyr- allele is recessive and results in albino homozygotes.  DNA sampling indicates that the frequency of tyr+ is 0.9 and tyr- is 0.1.  Pedigree studies indicate that the inbreeding coefficient is 0.3 in this population.  What proportion of the population should be albino?

How does this compare to expected frequency of albinos with the same allele frequencies but no inbreeding?

If the proportion of albinos is less than expected what might likely be occurring (assume that the estimate of inbreeding is correct).
Answer #1

Gene under consideration is- tyrosinase (tyr).

tyr+ is Dominant - expressed as normal pigment in homozygotes (tyr+, tyr+) and heterozygotes (tyr+, tyr-)

tyr- is Recessive - albino homozygotes (tyr-, try-)

Frequency of tyr+ = p =0.9

Frequency of tyr- = q =0.1

Note: p + q = 1, So population has Hardy-Weinberg equilibrium for allele frequency.

For population with inbreeding coefficient F =0.3, modified Hardy-Weinberg equilibrium will be

(p2 + pqF) + (2pq - 2pqF) + (q2 + pqF) =1

frequency of homozygotes recessive (tyr-, try-) albino in inbreeding population will be = (q2 + pqF)

(q2 + pqF) = (0.1 X 0.1) + (0.9 X 0.1 X 0.3) = 0.01 + 0.027 = 0.037

Thus 0.037 will be frequency of albino in inbreeding population

For population, Hardy-Weinberg equilibrium p2 + 2pq + q2 =1 assumes random mating (no inbreeding in population)

p2 is frequency of homozygotes dominant (tyr+, tyr+), 2pq is frequency of heterozygotes (tyr+, tyr-) and q2 is frequency of homozygotes recessive (tyr-, try-) albino

frequency of homozygotes recessive (tyr-, try-) albino will be = q2= 0.1 X 0.1 = 0.01

Expected frequency of albinos in population with random mating is 0.01

Thus inbreeding population will have higher frequency of albino in population.

If population is inbreeding, then higher frequency of albino in population is expected. However, if proportion of albino is less than expected, this could be due to higher mortality or premature death or other health deficiencies in albino gorillas as compared to heterozygotes and dominant homozygotes for tyr gene